working together, Dan and Leah can mow a lawn in 1hour and 12 min. working alone , Leah can mow the lawn in 1hour less time than it would take Dan. Find the time it takes for each of them working alone?
Answer by stanbon(75887)

working together, Dan and Leah can mow a lawn in 1hour and 12 min. working alone , Leah can mow the lawn in 1hour less time than it would take Dan. Find the time it takes for each of them working alone?
-----
Note: 1 hr 12 min = 6/5 hours
------
Dan DATA: time = x hrs/job ; rate = 1/x job/hr
----
Leah DATA: time = x-1 hrs/job ; rate = 1/(x-1) job/hr
-----
Together DATA: time = 6/5 hrs/job ; rate = 5/6 job/hr
----------------
Equation:
rate + rate = together rate
1/x + 1/(x-1) = 5/6
-----
6(x-1) + 6x = 5x(x-1)
-----
6x-6 + 6x = 5x^2 - 5x
5x^2 -17x +6 = 0
x = 3 hrs (Dan's time to do the job alone)
x-1 = 2 hrs (Leah's time to do the job alone)
-----
Cheers,
Stan H.
============